I need help finding the solution set for (x-2)(x-3)=2
heres my work: -2x-3x+6x^2-2=0
put them in order 6x^2-5x-2=0 A=6 B=-5 C=-2
5+-square root -5^2-4*6*-2/2*6
5+-SR 25+48/12
5+-SR 73/12
I am pretty sure that 73 is a prime number and don't know how to factor it and solve for x from here.


Thanks for what you do!

Answer by Alan3354(69443)

I need help finding the solution set for (x-2)(x-3)=2
----------------
x^2 - 5x + 6 = 2
x^2 - 5x + 4 = 0
(x-1)*(x-4) = 0
x = 1
x = 4
----------
heres my work: -2x-3x+6x^2-2=0 - You put the x^2 and the 6 together.
---------------------
put them in order 6x^2-5x-2=0 A=6 B=-5 C=-2
5+-square root -5^2-4*6*-2/2*6
5+-SR 25+48/12
5+-SR 73/12
I am pretty sure that 73 is a prime number and don't know how to factor it and solve for x from here.